OpenStudy (anonymous):

What are all the roots for r^4+1=0 and how do I find them?

OpenStudy (anonymous):

A root is anywhere where a function is equal to zero. Your equation equals zero when: \[r^4 = -1\] Which has no real roots.

OpenStudy (anonymous):

What are the non real roots

OpenStudy (anonymous):

Have you covered complex numbers? Are you expected to give the complex roots?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

\[r^4=\cos(\pi+2n \pi)+i \sin(\pi+2n \pi)\] since cos(pi)=-1 and sin(pi)=0 the right hand side is still -1 since -1+0=-1

myininaya (myininaya):

use demoirve's law (i think that is what it is called)

OpenStudy (anonymous):

The solution shows me radical2/2 +/- radical2/2 i; -rdicL2/2 =/- radical 2/2 i @myininaya but I'm not sure how to get those

myininaya (myininaya):

take the 4 root of both sides

OpenStudy (anonymous):

ohh right thank youuu!

myininaya (myininaya):

you would have one solution given by n=0 another by n=1 another by n=2 and the last by n=3

myininaya (myininaya):

that is 4 solutions